A project under consideration has an internal rate of return of 14% and a beta of 0.6. The risk-free rate is 9%, and the expected rate of return on the market portfolio is 14%.
a-1. Calculate the required return.
  Required return %  
a-2. Should the project be accepted?
   
 
  Yes
  No
b-1. Calculate the required return if its beta is 1.6.
  Required return %  
b-2. Should the project be accepted?
   
 
  Yes
  No
Answer #1

a-1) Required return = Risk-free rate + Beta * (Expected rate of return on market - Risk-free rate)

= 9 + 0.6 ( 14 - 9)

= 12%

a-2) Project should be accepted as the actual return of 14 % (Internal rate of return) is greater than the required return of 12%.

b-1) If New Beta is 1.6

Required return = Risk-free rate + Beta * (Expected rate of return on market - Risk-free rate)

= 9 + 1.6 ( 14 - 9)

= 17 %

b-2) Project should not be accepted as the actual return of 14 % (Internal rate of return) is less than the required return of 17 %.